Systems Staff Meeting Notes 3/9/98

Systems Staff Meeting Notes
March 9, 1998

Present: Rich Bennett, Brian Burton, Bill Covey, Destry Fudge, Debra Harris, Sharon Henry, Bobbie Parker, Suzy Shaw, Leota Whitfield

Announcements
  1. Weekend wiring went fine. There was only one network wire - to the RS6000 - that was flaky, & Suzy fixed that on Sunday.
  2. Workmen are scheduled to return to fix the panel box & finish caulking on Tuesday.
  3. Systems Meeting next Monday will be held in the Library East Conference Room with Martha Hruska in attendance.
  4. Students are all gone this week for Spring Break.
  5. Bobbie Parkerís title has been changed as of March 6 - she is now in the probationary period.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Switches still have not been received. They have been on order for 4 months now. They are the only solution to fixing Education and getting more IPís. Bill will call Scott at CISCO & try to track them down.
  2. CD ROM recordable drive for Preservation has not yet been received.
  3. Laptop external drive still not received.
  4. Stacey Ruszczyk is requesting more memory for a few machines in her area. Destry will get a head count & check the prices.
  5. John Ashcraft reported a problem with the 3rd floor of MSL - the Network Neighborhood wasnít visible. Destry fixed the problem - when the sharing is turned on, that machine becomes the master browser. He had to disable the master browser. Destry also notes that they need a new graphics-capable printer over there - HP6 or equivalent.
  6. The HP6 printer for Personnel has not yet been received.
  7. Maps - they are using a 6MP printer simply as a side printer. It has an internal network card already installed.
  8. Maps - they have a Color InkJet set up on a wobbly table that also houses a Cow. When the InkJet is printing, it really shakes that Gateway. They need to relocate the printer to a more stable location. Will speak with Tom Minton to pass the word along.
Networking
  1. Cable modem tests are being delayed.
  2. ADSL - from BellSouth - testing a new system. Must be within 20,000 cable feet (3 miles) of the big switch.
  3. Special Collections - most of the staff have been switched to the new wire, a few are still on the old due to IPís. Macs are on the new wiring, but AppleTalk is not yet set up on that port - so they canít print.
  4. North Tower & LAC is completely switched over. They are having problems with intermittent slow speeds.
  5. Education - 100 Mb fiber will be available soon.
  6. Systems/Preservation - have not heard back after submitting estimates.
  7. MSL - HVAC wiring - still no word.
  8. Library West Circulation Desk - now up & functional. There is one SNA terminal in the desk area. Tomorrow morning the IPís should work & the scanners are plugged in.
  9. The fire alarm in Circulation was inadvertently set off by hammering on some panduit that was close to it. The pull station needs to be checked, it is too sensitive.
  10. The baseboard electrical outlets on the cubicles in room 10 are now functional.
  11. The Resource Services router is in the same location - it has not been moved.
Software Projects Status
  1. Winston has spec for ILL migration to MS SQL out on a website for their review.
  2. Mac-Personnel switchover still in progress.
  3. It is unknown if Resource Services is finished with Mellon.
  4. Brian has completed the revision to the Systems Locator for the staff. He is currently working on the studentsí. The Locator is upgradable, it is table-driven.
  5. Brian is requesting more memory - says he has 16 & needs 32 RAM.
  6. Brian is also requesting Crystal Reports. Sharon has Version 6, Winston also needs an upgrade. They need to have three copies at the current level.
DP Coordinator
  1. Debra received a request from a Paul Zarins from Stanford University in California (request routed through Documents) for copies of National Survey of Hospital Data Processing. There is one 3480 cartridge & three 5 ľ floppy disks. LeiLani has located the hard copy material, says she can send it through ILL.
  2. The best way of transferring this information would be for him to set up a FTP account for us to upload the to. Debra will email him and try to work out the best way to get him the requested information.
Calendar
  1. Winston Harris out today.
  2. Spring Break March 9 - 13.
  3. St. Pattyís Day March 17.
  4. Destry out March 18 & 23.
  5. Shower for Marilyn & Destry - March 19, 11:30 AM - West Lounge. Marilyn is registered at Target & Toys ĎRí Us.
  6. Bobbie Parker out March 23 - 25.

    Back to the Systems Homepage

    Last updated March 7, 2000
    by Debra Harris
    debharr@mail.uflib.ufl.edu